OpenStudy (anonymous):

\frac{\frac{d^{2}v}{dx^{2}}}{[1+(\frac{dv}{dx})^{2}]^{\frac{3}{2}}}=cM(x) Find v=?

OpenStudy (jamesj):

What is this problem, physically?

OpenStudy (jamesj):

\[\frac{\frac{d^{2}v}{dx^{2}}}{[1+(\frac{dv}{dx})^{2}]^{\frac{3}{2}}}=cM(x)\]

OpenStudy (anonymous):

It's equation of deflection of beam

OpenStudy (jamesj):

Oh wait. Write w = dv/dx. Then this equation is w' / (1 + w^2)^(3/2) = = c M(x)

OpenStudy (jamesj):

and this equation is separable: \[\int\limits \frac{dw}{ (1+w^2)^{3/2}} \ = \ \int\limits cM(x) dx\]

OpenStudy (jamesj):

Now integrate, solve for w, then for v.

OpenStudy (anonymous):

after integrating \[\frac{v'}{\sqrt{1+(v')^{2}}+C_1=c\int M(x)dx\]

OpenStudy (anonymous):

\[\frac{v'}{\sqrt{1+(v')^{2}}}+C_1=c\int M(x)dx\] still not solved

OpenStudy (jamesj):

But once you have, you'll find the expression in w is quite messy and you'll be all but convinced that it's very difficult to write down a general solution for v. Which is why you need to go back to the problem; try and figure out more about v and the function M(x).

OpenStudy (jamesj):

Oh, sorry, yes you have written it down correctly. My mistake. But even so, solving for v' even now is still not pretty at all.

OpenStudy (anonymous):

put w=tan(@) [\dw=sec^{2}(@) d@\] [\int cos(@) d@\] so =sin@ =w/sqrt(1+w^2) +C1 since w=v' =v'/sqrt(1+(v')^2) +C1

OpenStudy (jamesj):

I..e, you can go through the algebra and find an expression for v' and then say v is the integral of that expression. But it's not going to be simple.

OpenStudy (anonymous):

I know it's a non linear equation and it's not simple. But I just want to find out it's solution.

OpenStudy (jamesj):

I.e., \[\frac{ v'^2 }{1 +v'^2} = ( c \int\limits M(x) dx - c_1)^2\] and now solve for v'.

OpenStudy (jamesj):

Well, keep going then. You can write down the expression.

OpenStudy (anonymous):

If I suppose v as power series then obviously the problem will not be solved

OpenStudy (jamesj):

No, I don't see how with a general M(x) power series help you.

OpenStudy (anonymous):

I found one Let \[f(x)=c\int M(x) dx-C_1\] \[v'^{2}=(1+(v')^{2})f(x)^{2}\] \[v'^{2}(1-f(x)^{2})=f(x)^{2}\] \[v'^{2}=\frac{f(x)^{2}}{1-f(x)^{2}}\] \[v'=\sqrt{\frac{f(x)^{2}}{1-f(x)^{2}}}\] so \[v=\int\sqrt{\frac{f(x)^{2}}{1-f(x)^{2}}}dx+C_2\]

OpenStudy (anonymous):

Thanks for your sincere help.